# 3D Foundation AI Model for Generalizable Disease Detection in Head Computed Tomography

### v1

@@ Results
To gain insight into the features learned through self-supervised pre-training and supervised fine-tuning of the foundation model, we visualize the attention maps within the Vision Transformer (ViT), as shown in Fig. 4. These heatmaps highlight the regions where the ViT model focuses most strongly. In the second column, we see that the pre-trained foundation model captures generic brain features, with dark red indicating attention on abnormal ventricular shapes and green marking areas of hemorrhage. After fine-tuning on specific tasks, the ViT’s attention becomes more focused on patterns relevant to each disease. For instance, in the edema task (third column), the heatmap extends across most of the brain, reflecting generalized swelling. For ADRD (fourth column), the model emphasizes regions of ventricular enlargement and cerebral atrophy. Multiple hemorrhages are also present in this sample, with attention covering both the IPH in the dense central region (fourth column) and extending toward the left end of the ventricle where IVH appears (fifth column). In the case of SAH (sixth column), the attention map is less prominent due to the small, peripheral area of the SAH in the lower part of the slice, although the model still predicts it accurately.
The comparison between the pre-trained and fine-tuned ViT explains the performance difference between linear probing and fine-tuning (shown in Supplementary Fig. 9, as end-to-end fine-tuning allows the model to learn task-specific features more effectively. Details on the computation of the visualized attention maps are provided in Section “Methods”.
@@ Discussion
Our study demonstrates that this pre-trained foundation model significantly outperforms models trained from scratch and other CT foundation models on the same labeled data. Moreover, it exhibits strong performance even with limited data, as shown in few-shot learning experiments, and suggests promising potential for scaling up with larger datasets. In clinical practice, head CT scans are typically acquired using heterogeneous protocols, including variations in slice thickness and scanner modalities. A robust foundation model for CT should generalize effectively across these diverse protocols. In this study, we utilized scans with slice thicknesses ranging from 0.5 mm to 5 mm and data from two major manufacturers (Siemens and Toshiba) to develop and assess the generalization capabilities of our foundation model. As illustrated in Supplementary Fig. 11, the embeddings produced by the foundation model show separability based on scanner manufacturer and slice thickness, likely reflecting variations in protocol distribution. However, by comparing the distribution of “All” patients to that of positive cases for each condition, we observe that the embeddings do not tend to collapse or bias towards a certain protocol. Supplementary Fig. 12 further demonstrates that fine-tuned models achieve comparable performance across scanner protocols. Detailed per-task performance results are provided in Supplementary Figs. 13 and 14. Additionally, in comparisons of Toshiba and Siemens scanners, we noted a systemically higher prevalence of positive cases across all tasks in Toshiba scans, leading to a modestly higher AUC in specific instances. Despite these variations, our foundation model demonstrates robust generalization capabilities across diverse CT protocols, highlighting its potential for broad clinical application.
@@ Methods
We utilized a large-scale head CT scan dataset from NYU Langone, consisting of 499,084 scans across 203,665 patients, collected between 2009 and 2023. These scans were acquired using Siemens and Toshiba machines. We included all the non-contrast head CT scans with ranging from 0.5mm to 5mm, kVp values between 70 and 150, and convolution kernels Hr/Qr/J with sharpness levels of 35-45. We filtered out corrupted scan series with missing DICOM files and those containing less than 10 slices, resulting in 451,298 scans. We partitioned these scans by the patient IDs into training, validation, and held-out validation sets in an 8:1:1 ratio to avoid the leakage of scans from the sample patient. As illustrated in Fig. 1a, this led to training, validation and held-out validation set with 361,663, 44,886 and 44,749 scans, respectively. The scans in the training set were used to train the foundation model.
@@
We evaluated our model using four datasets: one in-domain (ID) dataset from NYU Langone and three out-of-domain (OOD) datasets from NYU Long Island, the RSNA Challenge, and the public CQ500 dataset. Each dataset includes multiple head CT disease detection classes, with some classes abbreviated as follows: Hydrocephalus (HCP), Dementia (ADRD), Intraparenchymal Hemorrhage (IPH), Intraventricular Hemorrhage (IVH), Subdural Hemorrhage (SDH), Epidural Hemorrhage (EDH), Subarachnoid Hemorrhage (SAH), and Intracerebral Hemorrhage (ICH). These classes can have co-occur in the same Head CT scan. The characteristics of the patients are shown in Table 2. We split all datasets by patients to avoid information leakage, Further dataset details of our dataset are provided below.
NYU Langone - 10 tasks
The NYU Langone main campus dataset serves as the internal ID dataset for downstream evaluation. As depicted in Fig. 1b, patient health status was derived from Electronic Health Records (EHR) within a 3-month window centered around the scan date, with conditions defined by ICD-10 diagnostic codes and medications, outlined in Supplementary Table 1. This cohort includes 270,205 scans from 66,801 patients with valid EHRs, covering 10 classes: Tumor, HCP, Edema, ADRD, IPH, IVH, SDH, EDH, SAH, and ICH. This cohort was partitioned following the same split used used during pre-training: matched patients within the training, validation, and held-out subsets of the SSL pre-training phase were assigned to the corresponding sets of the supervised fine-tuning phase. This resulted in 217,109, 26,609, and 26,487 scans in the training, validation and test set, respectively.
NYU Long Island - 10 tasks
@@
As illustrated in Fig. 1b, we labeled head CT scans from NYU Langone and Long Island Hospital using electronic health records (EHR). For each head CT, we retrieved an EHR snippet for the corresponding patient based on their Medical Record Number (MRN), starting from the time of the scan and covering a 90-day period. We then checked for the presence of any diagnosis codes (ICD-10 codes) and medication records, within this EHR snippet that matched the predefined definitions for each disease, allowing us to create binary labels for each condition. The complete list of ICD-10 codes and the medications used for disease definitions is provided in Supplementary Table 1.
@@
For the 3D input volume, instead of creating 196 patches of size $16\times 16$ from a $224\times 224$ image as in standard 2D ViT, we segment $96\times 96\times 96$ 3D volumes into 512 patches of size $12\times 12\times 12$ for ViT input. This customized patch design considers the trade-off between performance and computational cost. As shown in Supplementary Fig. 4, our model outperforms a version using 216 patches of size $16\times 16\times 16$, indicating that smaller, more numerous patches enhance model performance. This supports the importance of capturing fine-grained features in 3D medical imaging, consistent with prior findings [19, 42]. However, computational costs increase significantly with respect to $s$ ($s$ defined as patch size reducing factor), at a rate of $O(s^{6})$, due to the cubic growth of patch numbers in 3D and the quadratic growth in self-attention computation (Supplementary D). To balance performance with computational efficiency, we adopt 512 patches of $12\times 12\times 12$ as the optimal input size for ViT in our foundation model.
@@
We compared the performance on downstream tasks between two versions of foundation models pre-trained using DINO and MAE, as shown in Fig. 3 and Supplementary Figs. 6, 9 and 8. The results indicate that DINO consistently outperforms MAE across all datasets. Based on this finding, we selected the DINO-pre-trained model as our final foundation model.
@@
Since no prior foundation models have been specifically trained on 3D Head CT for direct comparison, we benchmark our model against Merlin [20] and Google CT Foundation model [17] to highlight the advantages of our domain-specific foundation model. Merlin is a 3D Abdomen CT foundation model pre-trained on vision-language pairs with contrastive learning [5] and ICD code prediction task, where 6+ million images from 15,331 CTs, 1.8+ million diagnostic ICD codes from EHR, and 6+ million tokens from radiology reports are used. Different from our model architecture, Merlin used ResNet-152 ($\sim 60.4$M Parameters) as vision model with reshaped image size of $224\times 224\times 160$. The performance comparison between our model and Merlin is shown in Supplementary Fig. 5, where our model shows substantial improvement across most datasets and diseases. Google CT Foundation model is trained on a comprehensive private dataset comprising 527,078 CT studies with associated radiology reports from 430,772 patients. The model is first trained by Contrastive Captioning with CoCa [65] on 2D medical images and then adapting to CT by training on series of CT slices with VideoCoCa [66]. The performance comparison between our model and Google CT Foundation model is shown in Supplementary Fig. 7, where our model shows a consistent improvement across the board. We additionally show comparison of our model against model trained from scratch in Fig. 1 and Supplementary Fig. 5, where the overall significantly improved performance shows the effectiveness of our pre-training strategies on 3D Head CT images.
@@
We assessed pre-trained model performance through full fine-tuning (updating all weights) and various probing methods (updating only the classification layers). For both approaches, images were normalized to isotropic spacing, transformed to three HU interval channels, and reshaped to $3\times 96\times 96\times 96$. The entire transformed 3D image was then input into the ViT model for feature extraction, followed by an additional classification layer for downstream tasks. Probing utilized two strategies: linear probing, which adds a linear layer atop the ViT backbone, and attentive probing, which incorporates an attention layer. Attentive probing is chosen since MAE does not use [CLS] token as the learning objective. Linear probing only relies on [CLS] token to perform classification and attentive probing explores the interaction among all tokens [67]. Given the imbalances of downstream task labels, we randomly sampled a balanced subset from the training set per epoch, consisting of 5,000 samples (when fine-tuning on the NYU Langone, NYU Long Island, and RSNA datasets), and 500 samples when fine-tuning on CQ500. We trained all methods using the AdamW [58] optimizer with a cosine learning rate scheduler, a learning rate of $1\times 10^{-5}$ for backbone and $1\times 10^{-3}$ for classification layers, cross-entropy loss, and a maximum of 10 epochs. The main evaluation result with linear probing is shown in Fig. 1 with fine-tuning and probing comparison shown in Supplementary Fig. 6 for average performance across all diseases and Supplementary Fig. 9 for per disease performance. The result indicates that probing achieves performance levels close to full fine-tuning, underscoring the high quality of learned representations in our model.
@@
In order to evaluate the effectiveness of our model under scare label conditions, we applied few-shots learning where each class is only sampled $K$-times. Specifically, we chose $K=8,16,32,64,128,256$, where the data is sampled such that positive and negative samples equal to $K$ for each disease. Few-shot training was performed using full fine-tuning with the same hyper-parameter settings. While we also attempted some other commonly used few-shots classification methods such as k-nearest neighbors (KNN), Simple Shots [68] and Prototypical Networks [69], we did not observe performance improvement on our datasets over full fine-tuning. The main evaluation for few-shots classification is present in Fig. 2, where we observed our model can already reach performance close to full fine-tuning with only $K=256$ samples. This demonstrates the effectiveness of our model under scare data training regime.
@@
We visualize the average attention distances across all heads and layers for every patch in the volume in Fig. 4. This “attention distance” serves as an estimate of the ViT’s receptive field within the CT volumes, indicating the regions of the brain that the model focuses on. This visualization helps illustrate how the model integrates information across spatial areas to capture meaningful patterns within the volume.
@@
All experiments are performed under Python (v3.8.11), PyTorch (v2.4.1), CUDA (12.1) and MONAI (v1.2.0). We extend ViT, MAE, DINO implementation from original their corresponding repositories (https://github.com/facebookresearch/mae, https://github.com/facebookresearch/dino) to match our need for 3D CT image encoding. For comparison with Merlin [20], we integrated their original model weight checkpoints and model backbone code (https://github.com/louisblankemeier/merlin) to our downstream fine-tuning code base. ResNet50-3D [70] from (https://github.com/kenshohara/3D-ResNets-PyTorch/tree/master) is integrated to our code base for evaluation. All plots and figures were created by Matplotlib (v0.1.6) and Seaborn (v0.13.2). All downstream experiments were conducted on single 80 GB NVIDIA A100 GPU (graphics processing unit). All pre-training experiments were conducted on four 80 GB NVIDIA A100 GPUs.
@@ Appendix A Details of datasets
This section provides additional details about the dataset used to evaluate the downstream tasks. Table 1 lists the ICD-10 codes and medications used to identify the diagnoses for each disease. Table 2 presents the distribution of patient characteristics for each disease. Figs. 1 and 2 illustrates the prevalence of each disease in the downstream tasks for the NYU Langone and NYU Long Island datasets, highlighting the imbalances present in these tasks.
@@
Supplementary Table 2 Continued. Demographic characteristics of patients associated with scans from the NYU Long Island dataset, matched with electronic health records (EHR) and utilized in downstream tasks.

### v2

@@ Results
To gain insight into the features learned through self-supervised pre-training and supervised fine-tuning of the foundation model, we visualize the attention maps within the Vision Transformer (ViT), as shown in Fig.˜5. These heatmaps highlight the regions where the ViT model focuses most strongly. In the second column, we see that the pre-trained foundation model captures generic brain features, with dark red indicating attention on abnormal ventricular shapes and green marking areas of hemorrhage. After fine-tuning on specific tasks, the ViT’s attention becomes more focused on patterns relevant to each disease. For instance, in the edema task (third column), the heatmap extends across most of the brain, reflecting generalized swelling. For ADRD (fourth column), the model emphasizes regions of ventricular enlargement and cerebral atrophy. Multiple hemorrhages are also present in this sample, with attention covering both the IPH in the dense central region (fifth column) and extending toward the left end of the ventricle where IVH appears (sixth column). In the case of SAH (seventh column), the attention map is less prominent due to the small, peripheral area of the SAH in the lower part of the slice, although the model still predicts it accurately.
The comparison between the pre-trained and fine-tuned ViT explains the performance difference between linear probing and fine-tuning (shown in Supplementary Fig.˜14, as end-to-end fine-tuning allows the model to learn task-specific features more effectively. Details on the computation of the visualized attention maps are provided in Section “Methods”.
@@ Discussion
Our study demonstrates that this pre-trained foundation model significantly outperforms models trained from scratch and other CT foundation models on the same labeled data. Moreover, it exhibits strong performance even with limited data, as shown in few-shot learning experiments, and suggests promising potential for scaling up with larger datasets. In clinical practice, head CT scans are typically acquired using heterogeneous protocols, including variations in slice thickness and scanner modalities. A robust foundation model for CT should generalize effectively across these diverse protocols. In this study, we utilized scans with slice thicknesses ranging from 0.5 mm to 5 mm and data from two major manufacturers (Siemens and Toshiba) to develop and assess the generalization capabilities of our foundation model. As illustrated in Supplementary Fig.˜16, the embeddings produced by the foundation model show separability based on scanner manufacturer and slice thickness, likely reflecting variations in protocol distribution. However, by comparing the distribution of “All” patients to that of positive cases for each condition, we observe that the embeddings do not tend to collapse or bias towards a certain protocol. Supplementary Fig.˜17 further demonstrates that fine-tuned models achieve comparable performance across scanner protocols. Detailed per-task performance results are provided in Supplementary Figs.˜18 and 19. Additionally, in comparisons of Toshiba and Siemens scanners, we noted a systemically higher prevalence of positive cases across all tasks in Toshiba scans, leading to a modestly higher AUC in specific instances. Despite these variations, our foundation model demonstrates robust generalization capabilities across diverse CT protocols, highlighting its potential for broad clinical application.
@@ Methods
We utilized a large-scale head CT scan dataset from NYU Langone, consisting of 499,084 scans across 203,665 patients, collected between 2009 and 2023. These scans were acquired using Siemens and Toshiba machines. We included all the non-contrast head CT scans with ranging from 0.5mm to 5mm, kVp values between 70 and 150, and convolution kernels Hr/Qr/J with sharpness levels of 35-45. We filtered out corrupted scan series with missing DICOM files and those containing less than 10 slices, resulting in 451,298 scans. We partitioned these scans by the patient IDs into training, validation, and held-out validation sets in an 8:1:1 ratio to avoid the leakage of scans from the sample patient. As illustrated in Fig.˜1a, this led to training, validation and held-out validation set with 361,663, 44,886 and 44,749 scans, respectively. The scans in the training set were used to train the foundation model.
@@
We evaluated our model using four datasets: one in-domain (ID) dataset from NYU Langone and three out-of-domain (OOD) datasets from NYU Long Island, the RSNA Challenge, and the public CQ500 dataset. Each dataset includes multiple head CT disease detection classes, with some classes abbreviated as follows: Hydrocephalus (HCP), Dementia (ADRD), Intraparenchymal Hemorrhage (IPH), Intraventricular Hemorrhage (IVH), Subdural Hemorrhage (SDH), Epidural Hemorrhage (EDH), Subarachnoid Hemorrhage (SAH), and Intracerebral Hemorrhage (ICH). These classes can have co-occur in the same Head CT scan. The characteristics of the patients are shown in Table˜2. We split all datasets by patients to avoid information leakage, Further dataset details of our dataset are provided below.
NYU Langone - 10 detection tasks
The NYU Langone main campus dataset serves as the internal ID dataset for downstream evaluation. As depicted in Fig.˜1b, patient health status was derived from Electronic Health Records (EHR) within a 3-month window centered around the scan date, with conditions defined by ICD-10 diagnostic codes and medications, outlined in Supplementary Table˜1. This cohort includes 270,205 scans from 66,801 patients with valid EHRs, covering 10 classes: Tumor, HCP, Edema, ADRD, IPH, IVH, SDH, EDH, SAH, and ICH. This cohort was partitioned following the same split used used during pre-training: matched patients within the training, validation, and held-out subsets of the SSL pre-training phase were assigned to the corresponding sets of the supervised fine-tuning phase. This resulted in 217,109, 26,609, and 26,487 scans in the training, validation and test set, respectively.
NYU Long Island - 10 detection tasks
@@
RSNA — 4 retrieval tasks; CQ500 — 6 retrieval tasks
We evaluate volume-to-volume hemorrhage subtype retrieval on both RSNA and CQ500. For each positive case of a given subtype (IPH, IVH, SAH, and SDH in RSNA; ICH, IPH, IVH, SDH, EDH, and SAH in CQ500), the objective is to retrieve other cases with the same subtype. Retrieval is conducted in an all-vs-all setting, where every positive sample in the dataset is used once as a query against the remaining cases in the gallery (rest of samples in the dataset).
@@
As illustrated in Fig.˜1b, we labeled head CT scans from NYU Langone and Long Island Hospital using electronic health records (EHR). For each head CT, we retrieved an EHR snippet for the corresponding patient based on their Medical Record Number (MRN), starting from the time of the scan and covering a 90-day period. We then checked for the presence of any diagnosis codes (ICD-10 codes) and medication records, within this EHR snippet that matched the predefined definitions for each disease, allowing us to create binary labels for each condition. The complete list of ICD-10 codes and the medications used for disease definitions is provided in Supplementary Table˜1.
@@
For the 3D input volume, instead of creating 196 patches of size $16\times 16$ from a $224\times 224$ image as in standard 2D ViT, we segment $96\times 96\times 96$ 3D volumes into 512 patches of size $12\times 12\times 12$ for ViT input. This customized patch design considers the trade-off between performance and computational cost. As shown in Supplementary Fig.˜4, our model outperforms a version using 216 patches of size $16\times 16\times 16$, indicating that smaller, more numerous patches enhance model performance. This supports the importance of capturing fine-grained features in 3D medical imaging, consistent with prior findings [19, 45]. However, computational costs increase significantly with respect to $s$ ($s$ defined as patch size reducing factor), at a rate of $O(s^{6})$, due to the cubic growth of patch numbers in 3D and the quadratic growth in self-attention computation (Supplementary˜E). To balance performance with computational efficiency, we adopt 512 patches of $12\times 12\times 12$ as the optimal input size for ViT in our foundation model.
@@
We compared the performance on downstream tasks between two versions of foundation models pre-trained using DINO and MAE, as shown in Fig.˜4 and Supplementary Figs.˜6, 14 and 13. The results indicate that DINO consistently outperforms MAE across all datasets. Based on this finding, we selected the DINO-pre-trained model as our final foundation model.
@@
Since no prior foundation models have been specifically trained on 3D Head CT for direct comparison, we benchmark our model against Merlin [20] and Google CT Foundation model [17] to highlight the advantages of our domain-specific foundation model. Merlin is a 3D Abdomen CT foundation model pre-trained on vision-language pairs with contrastive learning [5] and ICD code prediction task, where 6+ million images from 15,331 CTs, 1.8+ million diagnostic ICD codes from EHR, and 6+ million tokens from radiology reports are used. Different from our model architecture, Merlin used ResNet-152 ($\sim 60.4$M Parameters) as vision model with reshaped image size of $224\times 224\times 160$. The performance comparison between our model and Merlin is shown in Supplementary Fig.˜5, where our model shows substantial improvement across most datasets and diseases. Google CT Foundation model is trained on a comprehensive private dataset comprising 527,078 CT studies with associated radiology reports from 430,772 patients. The model is first trained by Contrastive Captioning with CoCa [68] on 2D medical images and then adapting to CT by training on series of CT slices with VideoCoCa [69]. The performance comparison between our model and Google CT Foundation model is shown in Supplementary Fig.˜7, where our model shows a consistent improvement across the board. We additionally show comparison of our model against model trained from scratch in Fig.˜1 and Supplementary Fig.˜5, where the overall significantly improved performance shows the effectiveness of our pre-training strategies on 3D Head CT images.
@@
We assessed pre-trained model performance through full fine-tuning (updating all weights) and various probing methods (updating only the classification layers). For both approaches, images were normalized to isotropic spacing, transformed to three HU interval channels, and reshaped to $3\times 96\times 96\times 96$. The entire transformed 3D image was then input into the ViT model for feature extraction, followed by an additional classification layer for downstream tasks. Probing utilized two strategies: linear probing, which adds a linear layer atop the ViT backbone, and attentive probing, which incorporates an attention layer. Attentive probing is chosen since MAE does not use [CLS] token as the learning objective. Linear probing only relies on [CLS] token to perform classification and attentive probing explores the interaction among all tokens [70]. Given the imbalances of downstream task labels, we randomly sampled a balanced subset from the training set per epoch, consisting of 5,000 samples (when fine-tuning on the NYU Langone, NYU Long Island, and RSNA datasets), and 500 samples when fine-tuning on CQ500. We trained all methods using the AdamW [61] optimizer with a cosine learning rate scheduler, a learning rate of $1\times 10^{-5}$ for backbone and $1\times 10^{-3}$ for classification layers, cross-entropy loss, and a maximum of 10 epochs. The main evaluation result with linear probing is shown in Fig.˜1 with fine-tuning and probing comparison shown in Supplementary Fig.˜6 for average performance across all diseases and Supplementary Fig.˜14 for per disease performance. The result indicates that probing achieves performance levels close to full fine-tuning, underscoring the high quality of learned representations in our model.
@@
In order to evaluate the effectiveness of our model under scare label conditions, we applied few-shots learning where each class is only sampled $K$-times. Specifically, we chose $K=8,16,32,64,128,256$, where the data is sampled such that positive and negative samples equal to $K$ for each disease. Few-shot training was performed using full fine-tuning with the same hyper-parameter settings. While we also attempted some other commonly used few-shots classification methods such as k-nearest neighbors (KNN), Simple Shots [71] and Prototypical Networks [72], we did not observe performance improvement on our datasets over full fine-tuning. The main evaluation for few-shots classification is present in Fig.˜2, where we observed our model can already reach performance close to full fine-tuning with only $K=256$ samples. This demonstrates the effectiveness of our model under scare data training regime.
Volume-to-volume retrieval evaluation
We evaluate volume-to-volume hemorrhage subtype retrieval in an all vs. all setting. Each CT volume is represented by a fixed-length embedding stored along with multi-hot subtype labels. At evaluation time, embeddings are first $\ell 2$-normalized and a nearest-neighbor index is built on the gallery embeddings with inner-product search (equivalent to cosine similarity). For each subtype $c$, every positive case for $c$ is used once as a query and is retrieved against the entire gallery minus the query itself (self-match excluded). Relevance is defined as "same subtype cc". The ranked hit list per query yields a boolean relevance vector for which we compute mean Average Precision (mAP) and Precision@K@K (for $K=\{1,5,10\}$). Queries and gallery include all available positives (all-vs-all) for the evaluated datasets.
if a query has R relevant items in the galaeery, and the ranked list retrieves them at position $K=\{k_{1},k_{2},...,k_{R}\}$ ($K=\{1,3,5\}$ in our study), average precision for retreival is defined as $Ap=\frac{1}{R}\sum^{R}_{i=1}P(k_{i})$, where $P(k_{i})$ is the precision at rank $k_{i}$. Mean Average Precision is then defined as $mAP=\frac{1}{Q}\sum^{Q}_{q=1}AP(q)$ for query q and total number of queries Q.
@@
We visualize the average attention distances across all heads and layers for every patch in the volume in Fig.˜5. This “attention distance” serves as an estimate of the ViT’s receptive field within the CT volumes, indicating the regions of the brain that the model focuses on. This visualization helps illustrate how the model integrates information across spatial areas to capture meaningful patterns within the volume.
@@
All experiments are performed under Python (v3.8.11), PyTorch (v2.4.1), CUDA (12.1) and MONAI (v1.2.0). We extend ViT, MAE, DINO implementation from original their corresponding repositories (https://github.com/facebookresearch/mae, https://github.com/facebookresearch/dino) to match our need for 3D CT image encoding. For comparison with Merlin [20], we integrated their original model weight checkpoints and model backbone code (https://github.com/louisblankemeier/merlin) to our downstream fine-tuning code base. ResNet50-3D [73] from (https://github.com/kenshohara/3D-ResNets-PyTorch/tree/master) is integrated to our code base for evaluation. Nearest-neighbor indexing for volume-to-volume retrieval was done with faiss (v.1.12.0) (https://github.com/facebookresearch/faiss). All plots and figures were created by Matplotlib (v0.1.6) and Seaborn (v0.13.2). All downstream experiments were conducted on single 80 GB NVIDIA A100 GPU (graphics processing unit). All pre-training experiments were conducted on four 80 GB NVIDIA A100 GPUs.
@@ Appendix A Details of datasets
This section provides additional details about the dataset used to evaluate the downstream tasks. Table˜1 lists the ICD-10 codes and medications used to identify the diagnoses for each disease. Table˜2 presents the distribution of patient characteristics for each disease.
Figs.˜1 and 2 illustrates the prevalence of each disease in the downstream tasks for the NYU Langone and NYU Long Island datasets, highlighting the imbalances present in these tasks.
LABEL:tab:label_noise presents label sensitivity analysis based on ground truth labels from one radiologist on 1000 samples (891 samples in total after filtering duplicates) in comparison to ICD code mapped labels. The result is reported with Cohen’s Kappa $\kappa$, sensitivity and specificity. While majority of diseases show reasonably consistent agreement ($\kappa>0.3$ for $7$ out of $10$ diseases), specificity (>0.7>0.7 for $10$ out of $10$ diseases) and sensitivity (>0.7>0.7 for $7$ out of $10$ diseases). Epidural hemorrhage (EDH), edema and hydrocephalus show low agreement (<0.3<0.3) and low sensitivity (<0.5<0.5). Although sesativity of these three classes are low, when available at research grade labeling (i.e EDH available in CQ500), the direct external evaluation of the model trained on EHR based labeled data, without any fine-tuning on CQ500, reflects very high performance (See Fig.˜1)g. Since research-grade data is unavailable for Edema and Hydrocephalus we are not able to further verify this but we anticipate that the model performance even trained on low sensitivity labels (which are high in specificity) leads to sufficiently strong models.
@@
Supplementary Table˜2 Continued. Demographic characteristics of patients associated with scans from the NYU Long Island dataset, matched with electronic health records (EHR) and utilized in downstream tasks.
@@
Supplementary Table 3: Per-disease label sensitivity analysis
disease prevalence Cohen’s $\kappa$ sensitivity specificity accuracy
EDH 0.0078 0.0134 0.1429 0.9265 0.9203
ICH 0.1111 0.3761 0.9394 0.7525 0.7733
IPH 0.1111 0.5413 0.7677 0.9028 0.8878
IVH 0.0606 0.4513 0.7963 0.9092 0.9024
SAH 0.0955 0.6035 0.9529 0.8983 0.9035
SDH 0.0976 0.5645 0.8161 0.9104 0.9012
dementia 0.1470 0.7283 0.8855 0.9303 0.9237
edema 0.1279 0.0756 0.2719 0.8198 0.7497
hydrocephalus 0.0988 0.1118 0.2273 0.8941 0.8283
tumor 0.0539 0.5966 0.7917 0.9573 0.9484